OpenStudy (anonymous):

By with method do I solve x^4-37x^2+36 = 0?

hartnn (hartnn):

put x^2=y then x^4 = y^2 then u get a quadratic in y, can u solve that quadratic ?

hartnn (hartnn):

got that^ ?

OpenStudy (anonymous):

(y-36)(y-1), thx!

hartnn (hartnn):

yes, now find x

OpenStudy (anonymous):

"+ or -" 36, "+ or -" 1, got it!

hartnn (hartnn):

ummm, it will be \(\pm\)6, \(\pm\)1